# Toward Ultra-Long-Horizon Agentic Science: Cognitive Accumulation for Machine Learning Engineering

### v1

@@ 4.1 Experiment Setup
Baselines. To provide a comprehensive comparison, we compare ML-Master 2.0 with both open-source and closed-source baselines. Open-source baselines includes OpenHands [wang2025openhands], MLAB [huang2023mlagentbench], AIDE [jiang2025aide], R&D-Agent [yang2025rdagent] and AIRA-dojo [toledo2025aira]. Closed-source baselines includes Leeroo, Thesis, MLE-STAR-Pro [nam2025mlestar], FM Agent [li2025fmagent], etc. Due to the expensive cost of running complete MLEBench, we use the results reported by MLE-Bench as baselines.
@@ 4.2 Main Results
Table 1: ML-Master 2.0 outperforms all baselines on the main evaluation dimensions (percentage of achieving any medals across different machine learning task complexity levels) defined by MLE-Bench. Valid, Median+, Silver+ and Gold indicate the percentage of submissions with valid, above median score, silver medal or better and gold medal. Results are averaged over 3 runs with different random seeds and reported as mean ± one standard error of the mean (SEM). * denotes a incomplete report and the results are computed by padding incomplete seeds with failing scores. Best performances are marked in bold.
@@
ML-Master 2.0 achieves state-of-the-art performance across MLE-Bench. As shown in Table˜1, ML-Master 2.0 achieves the highest average medal rate of 56.4% among all evaluated methods, significantly outperforming the previous best closed-source method with a relative improvement of 11.2%(increasing from 50.7% to 56.4%) and surpassing the previous leading open-source method with a significant 60.7% improvement (from 35.1% to 56.4%).
@@
ML-Master 2.0 exhibits exceptional robustness and high-level human capability. ML-Master 2.0 maintains an average valid submission rate of 95.6%, which is on par with other leading approaches. Notably, it outperforms the top 50% of human participants in 63.1% of tasks, which is also the highest proportion among all methods, indicating exceptional robustness and a high performance floor.

### v2

@@ 4.1 Experiment Setup
Baselines. To provide a comprehensive comparison, we compare ML-ACE with both proprietary LLM-based methods and open-source LLM-based methods. Proprietary LLM-based methods includes OpenHands [wang2025openhands], MLAB [huang2023mlagentbench], AIDE [jiang2025aide], R&D-Agent [yang2025rdagent], AIRA-dojo [toledo2025aira], FM Agent [li2025fmagent] and MLE-STAR [nam2025mlestar], etc. Open-source LLM-based methods includes ML-Master [liu2025ml] and AutoMLGen [du2025automlgen]. Due to the expensive cost of running complete MLE-Bench, we use the results reported by MLE-Bench as baselines.
@@ 4.2 Main Results
Table 1: ML-Master 2.0 outperforms all baselines on the main evaluation dimensions (percentage of achieving any medals across different machine learning task complexity levels) defined by MLE-Bench. Valid, Median+, Silver+ and Gold indicate the percentage of submissions with valid, above median score, silver medal or better and gold medal. Results are averaged over 3 runs with different random seeds and reported as mean $\pm$ one standard error of the mean (SEM). * denotes a incomplete report and the results are computed by padding incomplete seeds with failing scores. Best performances are marked in bold.
@@
ML-ACE achieves state-of-the-art performance across MLE-Bench. As shown in Table˜1, ML-ACE achieves the highest average medal rate of 56.4% among all evaluated methods, significantly outperforming the previous best proprietary LLM-based methods method with a relative improvement of 11.2%(increasing from 50.7% to 56.4%).
@@
ML-Master 2.0 exhibits exceptional robustness and high-level human capability. ML-Master 2.0 maintains an average valid submission rate of 95.6%, which is on par with other leading approaches. Notably, it outperforms 50% human participants in 63.1% of tasks, which is also the highest among all methods, indicating exceptional robustness and a high performance floor.